# Happy family of stable marriages

### 02 Apr 2026 v1

@@ Conclusion
A strictly convex function defined on a bounded open set $\text{int}(K)$, which approaches infinity on the boundary of the set, must attain a global minimum at a unique interior point. Therefore, there exists a unique point that minimizes the self-perimeter. ∎
@@ 3.1 Self-Perimeter of a Regular $k$-gon
For regular $k$-gons, as well as for the family of planar convex bodies possessing $2\pi/k$ rotational symmetry, the self-perimeter exhibits periodic behavior depending on the remainder of $k$ modulo four.
Let $K$ be a regular $k$-gon centered at the origin, with circumradius $R$. The Euclidean length of each edge is $L=2R\sin(\pi/k)$. By symmetry, the self-perimeter is $k$ times the contribution of a single edge:
@@
where $r$ is the distance from the origin to the boundary in the direction parallel to the edge.
Assuming a vertex is at angle 0, the first edge connects the vertices at angles 0 and $2\pi/k$. The vector parallel to this edge has the angle $\alpha=\frac{\pi}{2}+\frac{\pi}{k}$. We cast a ray from the origin at angle $\alpha$. The boundary of $K$ consists of line segments with normal vectors at angles $\phi_{m}=\frac{(2m+1)\pi}{k}$ for integer $m$. The perpendicular distance from the origin to any edge is $d=R\cos(\pi/k)$. The ray intersects the $m$-th edge at a distance:
@@
    The regular polygon constitutes a maximum (upper bound) for the family of shapes with the same degree of symmetry.
@@
3.2 Self-Perimeter of Non-Perfect $k$-gons and Triangles
For a non-centrally symmetric polygon and an interior reference point $p$, we define the self-perimeter $P$ by integrating along the boundary.
Since the direction is constant along each straight edge, the integral simplifies to a discrete sum over the edges:
$P=\sum_{i=1}^{k}\frac{L_{i}}{r_{i}}$
where $L_{i}$ is the Euclidean length of the $i$-th edge, and $r_{i}$ is the length of the segment from $p$ to the boundary in the direction of $E_{i}$.
For a triangle with Euclidean edge lengths $L_{1},L_{2},L_{3}$, the sum is $P=\frac{L_{1}}{r_{1}}+\frac{L_{2}}{r_{2}}+\frac{L_{3}}{r_{3}}$ Using barycentric coordinates $p=\alpha V_{1}+\beta V_{2}+\gamma V_{3}$ where $\alpha+\beta+\gamma=1$, the edge segment lengths evaluate to $r_{1}=\alpha L_{1}$, $r_{2}=\beta L_{2}$, and $r_{3}=\gamma L_{3}$. The total self-perimeter is therefore $P(\alpha,\beta,\gamma)=\frac{1}{\alpha}+\frac{1}{\beta}+\frac{1}{\gamma}$.
To find the minimum possible self-perimeter, this functional is minimized to achieve a minimum of 9, which occurs at the centroid where $\alpha=\beta=\gamma=\frac{1}{3}$.
Theorem 3.3 (Grunbaum[7]).
@@ 3.2 Self-Perimeter of Non-Perfect $k$-gons and Triangles
4 Self perimeter in $\mathbb{R}^{n}$
First, let us define the perimeter $P^{(n)}_{B}$ of a center symmetric convex set $B\subset\mathbb{R}^{n}$. Once we defined this perimeter, we postulate that the $n$- volume of $B$ is just $V_{B}^{(n)}=P^{(n)}_{B}/n$.
@@ 4 Self perimeter in $\mathbb{R}^{n}$
$nV(B^{(n)})=P(B^{(n)}):=\int_{\partial^{*}B}\frac{V(B^{(n-1)}(\nu_{x}))d{\cal H}_{n-1}(x)}{{\cal H}_{n-1}(B^{(n-1)}(\nu_{x}))}$ (5)
where $\partial^{*}B$ is the essential boundary of $B$ and $\nu_{x}$ is the normal to $\partial B$ defined everywhere on $x\in\partial^{*}B$ .
@@
Note that (5) is consistent with (2) provided $V(B^{(1)})=2$. Indeed, if $n=2$ then ${\cal H}_{1}(B^{(1)}(\phi))=2r_{B}(\phi)$ because $B$ is center symmetric.
4.1 Invariance of the Self-Perimeter Under Affine Transformations
In this section, we prove that the self-perimeter $P(B^{(n)})$, and consequently the self-volume $V(B^{(n)})=\frac{1}{n}P(B^{(n)})$, is invariant under any invertible affine transformation $T:\mathbb{R}^{n}\to\mathbb{R}^{n}$ that preserves the origin. We proceed by mathematical induction on the dimension $n$.
Base Case: $n=2$
Lemma 4.1.
Let $B\subset\mathbb{R}^{2}$ be a convex set containing the origin. For any invertible linear transformation $T:\mathbb{R}^{2}\to\mathbb{R}^{2}$, the self-perimeter is invariant: $P(T(B))=P(B)$.
@@ 4.1 Invariance of the Self-Perimeter Under Affine Transformations
Let the boundary $\partial B$ be parameterized by a curve $\gamma(t)$ for $t\in[a,b]$. The self-perimeter in $\mathbb{R}^{2}$ is defined as:
$P(B)=\int_{\partial B}\frac{ds}{r_{B}(\vec{\tau}(s))}$ (6)
where $ds=\|\gamma^{\prime}(t)\|dt$ is the Euclidean arc length element, and $\vec{\tau}(t)=\frac{\gamma^{\prime}(t)}{\|\gamma^{\prime}(t)\|}$ is the unit tangent vector.
The radius function $r_{B}(v)$ is homogeneous of degree 1-1 along rays, meaning for any scalar $c>0$, $r_{B}(cv)=\frac{1}{c}r_{B}(v)$. Applying this to the unit tangent vector yields:
$r_{B}(\vec{\tau}(t))=r_{B}\left(\frac{\gamma^{\prime}(t)}{\|\gamma^{\prime}(t)\|}\right)=\|\gamma^{\prime}(t)\|r_{B}(\gamma^{\prime}(t))$ (7)
Substituting this back into the integral, the Euclidean norm cancels out completely:
$P(B)=\int_{a}^{b}\frac{\|\gamma^{\prime}(t)\|dt}{\|\gamma^{\prime}(t)\|r_{B}(\gamma^{\prime}(t))}=\int_{a}^{b}\frac{dt}{r_{B}(\gamma^{\prime}(t))}$ (8)
Now consider the transformed body $\tilde{B}=T(B)$. Its boundary $\partial\tilde{B}$ is parameterized by $\tilde{\gamma}(t)=T(\gamma(t))$, which gives $\tilde{\gamma}^{\prime}(t)=T(\gamma^{\prime}(t))$. The radius function of the transformed body for a vector $T(v)$ is:
$r_{\tilde{B}}(T(v))=\sup\{\lambda>0:\lambda T(v)\in T(B)\}$ (9)
Since $T$ is linear and invertible, $\lambda T(v)\in T(B)$ if and only if $\lambda v\in B$. Therefore, $r_{\tilde{B}}(T(v))=r_{B}(v)$. This implies $r_{\tilde{B}}(\tilde{\gamma}^{\prime}(t))=r_{B}(\gamma^{\prime}(t))$.
Computing the self-perimeter for $\tilde{B}$ using the parameterization formula:
$P(\tilde{B})=\int_{a}^{b}\frac{dt}{r_{\tilde{B}}(\tilde{\gamma}^{\prime}(t))}=\int_{a}^{b}\frac{dt}{r_{\tilde{B}}(T(\gamma^{\prime}(t)))}=\int_{a}^{b}\frac{dt}{r_{B}(\gamma^{\prime}(t))}=P(B)$ (10)
Thus, the base case $n=2$ holds. ∎
Inductive Step: $\mathbb{R}^{n}$
Theorem 4.1.
If $V(B^{(k)})$ is invariant under origin-preserving affine transformations for all dimensions $k<n$, then $P(B^{(n)})$ and $V(B^{(n)})$ are invariant in dimension $n$.

### 09 Apr 2026 v2

@@ Conclusion
A strictly convex function defined on a bounded open set $\text{int}(K)$, which approaches infinity on the boundary of the set, must attain a global minimum at a unique interior point. Therefore, there exists a unique point that minimizes the self-perimeter.
@@ 3.1 Self-Perimeter of a Regular $k$-gon
For regular $k$-gons, as well as for the family of planar convex bodies possessing $2\pi/k$ rotational symmetry, the self-perimeter exhibits periodic behavior depending on the remainder of $k$ modulo four. Let $K$ be a regular $k$-gon centered at the origin, with circumradius $R$. The Euclidean length of each edge is $L=2R\sin(\pi/k)$. By symmetry, the self-perimeter is $k$ times the contribution of a single edge:
@@
where $r$ is the distance from the origin to the boundary in the direction parallel to the edge. Assuming a vertex is at angle 0, the first edge connects the vertices at angles 0 and $2\pi/k$. The vector parallel to this edge has the angle $\alpha=\frac{\pi}{2}+\frac{\pi}{k}$. We cast a ray from the origin at angle $\alpha$. The boundary of $K$ consists of line segments with normal vectors at angles $\phi_{m}=\frac{(2m+1)\pi}{k}$ for integer $m$. The perpendicular distance from the origin to any edge is $d=R\cos(\pi/k)$. The ray intersects the $m$-th edge at a distance:
@@
3.2 Busemann’s definition
In the case of non-centrally symmetric sets, Busemann’s definition deviates from (3). Busemann considered the integral:
$P_{Bus}(B)=\int_{\partial B}\frac{2ds}{r_{B}(\vec{\tau}(s))+r_{B}(-\vec{\tau}(s))}$ (5)
namely, the integral is taken over twice the inverse length of the chord parallel to the tangent and passing through the origin. Note that $P_{Bus}(B)=P(B)$ for $B\in CCS$ and that in general, unlike $P$, $P_{Bus}$ is independent of the orientation of $\partial B$ (Remark 3.1). This definition is more natural when extending to higher dimensions, where the length of the chord is replaced by the Lebesgue measure of the parallel hyperplane (see Sec 4). Evidently, for CCS sets $P_{Bus}=P$ since $r_{B}(\phi)=r_{B}(-\phi)$.
Theorem 3.2 is valid also for $P_{Bus}$. The proof of this theorem can be extended to $P_{Bus}$ since $v\rightarrow r_{K,p}(-v)$ and $v\rightarrow r_{K,p}(v)+r_{K,p}(-v)$ are concave as well. Surprisingly, we did not find an explicit formulation of this result in the literature, so we formulate it below:
Theorem $3.$
Let $K\subset\mathbb{R}^{2}$ be a bounded convex set and $p\in Int(K)$. Then $p\rightarrow P_{Bus}(K-p)$ is strongly convex in $Int(K)$. In particular, there exists a unique interior point $p\in\text{int}(K)$ that minimizes the Busemann self-perimeter $P_{Bus}(K-\{p\})$.
The following Theorem can also be easily extended to Busemann’s self-perimeter:
Theorem 3.4 (Grünbaum & Martini [7]).
@@ 3.2 Busemann’s definition
Example
Let $T\subset\mathbb{R}^{2}$ be a triangle with Euclidean edge lengths $L_{1},L_{2},L_{3}$ and corresponding vertices $V_{1},V_{2},V_{3}$. Let $p\in\text{int}(T)$ be an interior reference point.
The point $p$ can be uniquely represented by its strictly positive barycentric coordinates $(\alpha,\beta,\gamma)$ such that:
$p=\alpha V_{1}+\beta V_{2}+\gamma V_{3}\quad\text{where}\quad\alpha+\beta+\gamma=1$ (7)
The asymmetric self-perimeter (3) relies on the lengths of the single directional rays $r_{i}$ originating from $p$ to the boundary in the directions parallel to the edges. For a triangle, these ray lengths evaluate to $r_{1}=\alpha L_{1}$, $r_{2}=\beta L_{2}$, and $r_{3}=\gamma L_{3}$. The self-perimeter is the discrete sum over the edges:
$P_{ray}(p)=\sum_{i=1}^{3}\frac{L_{i}}{r_{i}}=\frac{L_{1}}{\alpha L_{1}}+\frac{L_{2}}{\beta L_{2}}+\frac{L_{3}}{\gamma L_{3}}$ (8)
Which simplifies entirely to a function of the barycentric coordinates:
$P(\alpha,\beta,\gamma)=\frac{1}{\alpha}+\frac{1}{\beta}+\frac{1}{\gamma}$ (9)
Note that this perimeter is independent of the orientation of the triangle since the opposite orientation just permutes the barycentric coordinates but preserves the sum.
The Busemann perimeter (5) normalizes the Euclidean surface element using the full length of the parallel chords $c_{i}$ passing through $p$, rather than single rays. The full chord length parallel to an edge is the sum of the forward ray and the backward ray: $c_{i}=r_{i}+r_{i}^{-}$. For a 2D triangle, the Busemann definition multiplies the sum by $\omega_{1}=2$. The chords parallel to the edges evaluate to $c_{1}=(1-\alpha)L_{1}$, $c_{2}=(1-\beta)L_{2}$, and $c_{3}=(1-\gamma)L_{3}$. The Busemann self-perimeter is given by:
$P_{Bus}(p)=\sum_{i=1}^{3}\frac{2L_{i}}{c_{i}}=\frac{2L_{1}}{(1-\alpha)L_{1}}+\frac{2L_{2}}{(1-\beta)L_{2}}+\frac{2L_{3}}{(1-\gamma)L_{3}}$ (10)
Which simplifies to:
$P_{Bus}(\alpha,\beta,\gamma)=2\left(\frac{1}{1-\alpha}+\frac{1}{1-\beta}+\frac{1}{1-\gamma}\right)$ (11)
Since $\alpha+\beta+\gamma=1$, we can write:
$P_{Bus}(\alpha,\beta,\gamma)=2\left(\frac{1}{\beta+\gamma}+\frac{1}{\alpha+\gamma}+\frac{1}{\alpha+\beta}\right)\ .$
By the inequality $a^{-1}+b^{-1}\geq 4(a+b)^{-1}$, we obtain:
$9\leq P_{Bus}(T)\leq P(T)$
where both equalities hold whenever $p$ is the centroid $\alpha=\beta=\gamma=1/3$.
4 Self-perimeter in $\mathbb{R}^{n}$
We define the perimeter $|\partial K|_{B}$ and the affine invariant perimeter $P_{B}(K)$ of a set $K$ in a normed space $(\mathbb{R}^{n},\|\cdot\|_{B})$ where $B\subset\mathbb{R}^{n}$ is a convex, centrally symmetric set, by:
$|\partial K|_{B}:=\int_{\partial K}\omega^{(n-1)}_{B}(\nu_{x})d{\cal H}_{n-1}(x)\ \ ;\ \ P_{B}(K):=\int_{\partial K}\frac{\omega^{(n-1)}_{B}(\nu_{x})}{{\cal H}_{n-1}(\Sigma^{n-1}(\nu_{x}))}d{\cal H}_{n-1}(x)$
where $\Sigma^{n-1}(\nu_{x})$ is the (n1)(n-1)-subspace perpendicular to the normal $\nu_{x}\in\partial K$, and $\omega_{B}^{(n-1)}$ is defined recursively via:
Definition 4.1.
Given an orthonormal frame $\nu_{k+1},\ldots\nu_{n}$ in $\mathbb{R}^{n}$, let $B^{(k)}(\nu_{k+1},\ldots,\nu_{n})\subset\mathbb{R}^{k}$ be the intersection of $B$ with the subspace orthogonal to $\text{Sp}(\nu_{k+1},\ldots,\nu_{n})$. Define:
$\omega_{B}^{(k)}(\nu_{k+1},\ldots\nu_{n})=k^{-1}\int_{\partial B^{(k)}(\nu_{k+1},\ldots,\nu_{n})}\frac{\omega_{B}^{(k-1)}(\nu_{x},\nu_{k+1},\ldots,\nu_{n})}{{\cal H}_{k-1}(B^{(k-1)}(\nu_{x},\nu_{k+1},\ldots,\nu_{n}))}d{\cal H}_{k-1}(x)$ (12)
for $k=2,\ldots,n-1$, where $\nu_{x}\in\text{Sp}^{\perp}(\nu_{k+1},\ldots,\nu_{n})$ is the normal to $\partial B^{(k)}(\nu_{k+1},\ldots,\nu_{n})$ at $x$. This recursive definition starts at $k=2$ with $\omega^{(1)}(\nu_{2},\ldots,\nu_{n})\equiv 2$, independent of the frame $\nu_{2},\ldots,\nu_{n}$.
In particular:
$\omega_{B}^{(n)}=n^{-1}\int_{\partial B^{(n)}}\frac{\omega^{(n-1)}_{B}(\nu_{x})d{\cal H}_{n-1}(x)}{{\cal H}_{n-1}(B^{(n-1)}(\nu_{x}))}$ (13)
@@ 4 Self-perimeter in $\mathbb{R}^{n}$
Note that for $n=2$, (13) is consistent with (3) since $\omega^{(1)}=2$. Indeed, if $n=2$, then ${\cal H}_{1}(B^{(1)}(\phi))=2r_{B}(\phi)$ because $B$ is centrally symmetric. However, if $B$ is not centrally symmetric, then (13) deviates from (3). Altogether, Definition 4.1 is valid for any convex set $B$ provided $\{0\}\in Int(B)$.
4.1 Examples
The hypercube $1,1]^{n}$
The central cross-section of $B_{n}=[-1,1]^{n}$ passing through the origin and orthogonal to its normal is $B_{n-1}=[-1,1]^{n-1}$. The Euclidean (n1)(n-1)-dimensional volume of this cross-section is $\mathcal{H}_{n-1}(B_{n-1})=2^{n-1}$. Because the normal vector is constant across a single face, the integrand is also constant. The integral over the boundary $\partial B_{n}$ is the sum of the integrals over the $2n$ faces:
$\omega^{(n)}_{B_{n}}=n^{-1}\sum_{k=1}^{2n}\int_{\text{face}_{k}}\frac{\omega^{(n-1)}_{B_{n-1}}(\nu_{k})}{2^{n-1}}d\mathcal{H}_{n-1}(x)$
where $\nu_{k}$ is the constant unit vector perpendicular to the $k$-th face. By Theorem 4.4, it follows that $\omega^{(n-1)}_{B_{n-1}}(\nu_{k})$ is independent of $k$, hence $\omega^{(n-1)}_{B_{n-1}}(\nu_{k})=\omega^{(n-1)}_{B_{n-1}}$. In addition, the Lebesgue measure of $B_{n-1}$ is $2^{n-1}$, so:
$\omega^{(n)}_{B_{n}}=n^{-1}\times 2n\times\omega^{(n-1)}_{B_{n-1}}=2\omega^{(n-1)}_{B_{n-1}}$
Since $\omega^{(1)}_{B_{1}}=2$, it follows:
$\omega^{(n)}_{B_{n}}=2^{n}$
in agreement with the Lebesgue measure of $1,1]^{n}$.
The Euclidean ball
Since the Lebesgue measures of the cross-sections of the Euclidean ball are all constants equal to $\omega^{(n-1)}$, it follows that the self-perimeter is:
$\omega_{B}^{(n)}=n^{-1}\int_{{\mathbb{S}^{n-1}}}\frac{\omega_{B}^{(n-1)}(\nu_{x})}{\omega^{(n-1)}}d{\cal H}_{n-1}(x)$
Since $\int_{{\mathbb{S}^{n-1}}}d{\cal H}_{n-1}=n\omega^{(n)}$ and $\omega_{B}^{(n-1)}$ is independent of orientation, we obtain:
$\omega_{B}^{(n)}=\frac{\omega_{B}^{(n-1)}\omega^{(n)}}{\omega^{(n-1)}}\ .$
Since $\omega^{(1)}=\omega_{B}^{(1)}=2$, we obtain the equality $\omega_{B}^{(n)}=\omega^{(n)}$ for any $n\geq 2$.
The $n$-simplex
In this example, we study the self-volume of a non-centrally symmetric ball. Let $\Delta^{n}$ be an $n$-dimensional simplex with $n+1$ vertices $v_{1},\ldots,v_{n+1}\in\mathbb{R}^{n}$. Any interior point $p\in\text{int}(\Delta^{n})$ can be uniquely represented by its strictly positive barycentric coordinates $\lambda=(\lambda_{1},\dots,\lambda_{n+1})$, where $p=\sum_{i=1}^{n+1}\lambda_{i}v_{i}$ and $\sum_{i=1}^{n+1}\lambda_{i}=1$, with $\lambda_{i}>0$ for all ii. Let $F_{i}$ denote the (n1)(n-1)-dimensional face opposite to the $i$-th vertex.
Theorem 4.1.
The self-volume of the simplex $\Delta^{(n)}$ with center at $p$ is:
$\omega_{\Delta^{n}}^{(n)}=\frac{2}{n!}\sum_{k_{1},\dots,k_{n-1}}\prod_{m=1}^{n-1}\frac{1}{1-\sum_{r=1}^{m}\lambda_{k_{r}}}$
where $\{k_{1},\ldots,k_{n-1}\}\subset\{1,\ldots,n+1\}$ are n1n-1 different indices ($k_{i}\neq k_{j}$ for $i\neq j$) and $\lambda_{1},\ldots,\lambda_{n+1}$ are the barycentric coordinates of pp.
@@ 4.1 Examples
By applying the recursive definition of $\omega_{\Delta^{n}}$ (12), the integral over the boundary $\partial\Delta^{n}$ decomposes into a sum over the $n+1$ faces. For a specific face $F_{i}$, the integration is evaluated over the (n1)(n-1)-dimensional cross-section $S_{i}$ passing through $p$ and parallel to $F_{i}$. Because $S_{i}$ is parallel to $F_{i}$, it is a simplex similar to $F_{i}$ with a linear scaling factor of $(1-\lambda_{i})$. Consequently, the ratio of their Euclidean volumes, which appears in the denominator, is given by:
Base Case ($n=1$):
For $n=1$, the sum of indices cancels out and the empty product is defined as 1. We obtain:
$\omega_{\Delta^{1}}^{(1)}=\frac{2}{1!}=2$
Inductive Step:
Assume the formula holds for $n$:
$\omega_{\Delta^{n}}^{(n)}=\frac{2}{n!}\sum_{k_{1},\dots,k_{n-1}}\prod_{m=1}^{n-1}\frac{1}{1-\sum_{r=1}^{m}\lambda_{k_{r}}}$
We will prove its validity for $n+1$. Based on the decomposition properties of the volume, the following recursive formula holds:
$\omega_{\Delta^{n+1}}^{(n+1)}=\frac{1}{n+1}\sum_{k_{n}}\frac{1}{1-\sum_{r=1}^{n}\lambda_{k_{r}}}\omega^{(n)}_{\Delta^{n}}$
Substitute the expression for $\omega_{\Delta^{n}}^{(n)}$ from the inductive hypothesis into the recursion:
$\omega_{\Delta^{n+1}}^{(n+1)}=\frac{1}{n+1}\sum_{k_{n}}\frac{1}{1-\sum_{r=1}^{n}\lambda_{k_{r}}}\left(\frac{2}{n!}\sum_{k_{1},\dots,k_{n-1}}\prod_{m=1}^{n-1}\frac{1}{1-\sum_{r=1}^{m}\lambda_{k_{r}}}\right)$
$\frac{2}{(n+1)!}\sum_{k_{1},\dots,k_{n}}\prod_{m=1}^{n}\frac{1}{1-\sum_{r=1}^{m}\lambda_{k_{r}}}\ .$
4.2 Convexity of self-volume
By Theorem 4.1 we obtain:
Corollary 4.1.
The self-volume $\omega^{(n)}(\Delta^{n}-p)$ is a strictly convex function of the center pp. In particular, there exists a unique $p^{*}\in\Delta^{n}$ for which $\omega^{(n)}(\Delta^{n}-p^{*})$ attains its minimum, and $\lambda_{i}=\frac{1}{n+1}$, $i=1,\ldots,n+1$ are the barycentric coordinates of $p^{*}$.
$\min_{p}\omega_{\Delta^{n}}^{(n)}=\frac{(n+1)^{n}}{n!}\ .$